# UAlign: Leveraging Uncertainty Estimations for Factuality Alignment on Large Language Models

### v1

@@ Appendix B Dataset Details
Appendix C Evaluation Details
Accuracy
@@ Appendix C Evaluation Details
Area Under the Receiver Operator Characteristic Curve (AUROC)
AUROC assesses the effectiveness of confidence estimation Filos et al. (2019) by quantifying how likely a randomly chosen correct answer possesses a higher confidence score than an incorrect one, yielding a score within the range of [0, 1], implemented by sklearn toolkit  [^8] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py. A higher AUROC score implying higher reliability is preferred.
Appendix D Baseline Details
@@
For all in-context learning methods, we extract the examples from the respective training set to mitigate the knowledge distribution shift between different datasets. For example, the demonstrated examples in Appendix I are derived from the TriviaQA training set and are specifically used when inferring on the TriviaQA validation set. For LSQA without the training set, we use the same examples as TriviaQA as their knowledge domains largely overlap.
  • ICL: Few-shot prompts containing $m$ examples are utilized for answer generation with temperature $T=0.2$ where $m$ is set to 2 as presented in the Template E.
  • ICL-IDK: Two examples are included in the few-shot prompt while one is selected from the ICL-used example, and another is an unknown question whose answer is revised to “Sorry, I don’t know.” as presented in the Template E.
  • ICL-CoT: We also employ the Chain-of-Thought in few-shot examples by recalling the relevant knowledge piece of LLMs and incorporating it into thinking steps before answering the question as presented in the Template E.
@@
Figure 7: Illustration of several baselines as in Sec. 3.3.
Appendix E Prompt Template
@@
Appendix F Training Setting Details
@@
Appendix G Detailed Related Works
G.1 Knowledge Boundary
Previous works investigate the knowledge boundary to identify the known level of a knowledge piece of LLMs by quantifying the confidence or uncertainty estimations like output consistency (Cheng et al., 2024), prompting methods (Ren et al., 2023) or knowledge probing (Ji et al., 2024). Researchers are examining the limits of parametric knowledge in LLMs with the objective of delineating the extent of the LLMs’ knowledge and identifying their capability boundaries. Present studies on the knowledge boundary primarily focus on measuring the knowledge boundaries using confidence or uncertainty estimations on specialized tasks. The ambiguity of knowledge boundaries can be attributed to the knowledge distribution learned from the pre-training stage or the influence of external knowledge leading to knowledge conflict (Xu et al., 2024b) and inconsistency (Xue et al., 2023).
@@ G.1 Knowledge Boundary
G.2 Uncertainty Estimation of LLMs
To alleviate over-confidence and enhance the reliability of LLMs, reliable uncertainty estimation is essential to determine whether a question is known or not to the LLM (Geng et al., 2023). Both Uncertainty and Confidence estimations can indicate the reliability degree of the responses generated by LLMs, and are generally used interchangeably Xiao et al. (2022); Chen and Mueller (2023); Geng et al. (2023); Lu et al. (2024). In this part, we investigate several commonly used confidence & uncertainty estimation methods for generative LLMs as mentioned in Sec. 5. Specifically, we denote $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ as the confidence score associated with the output sequence $\boldsymbol{y}=[y_{1},y_{2},\dots,y_{N}]$ given the input context $\boldsymbol{x}=[x_{1},x_{2},\dots,x_{M}]$. We also illustrate the summarized estimation methods as well as their disadvantages in Fig. 8.
@@ G.2 Uncertainty Estimation of LLMs
G.3 Factuality Alignment of LLMs
Alignment is a standard procedure to improve LLMs’ helpfulness and factuality (Bai et al., 2022a). The main goal of LLM alignment is to guide human preference through Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022a) or AI feedback (Bai et al., 2022b), which may also guide LLMs to output detailed and lengthy responses (Singhal et al., 2023) but inevitably encourage hallucination. Therefore, many works explore to apply RL to improve LLMs’ factuality through Proximal Policy Optimization (PPO) (Schulman et al., 2017) with the trained reward model (Liang et al., 2024; Xu et al., 2024a) or Direct Preference Optimization (DPO) Rafailov et al. (2023) with the constructed preference dataset (Zhang et al., 2024b; Lin et al., 2024) to align with factuality preferences annotated by human beings. Xu et al. (2024a) encourage LLM to reject unknown questions using the constructed preference data by leveraging knowledge boundary feedback.
Appendix H Experiments
H.1 Experiments of Reliability of Uncertainty Estimations
@@
Appendix I Few-shot Prompt Examples

### v2

@@ Appendix C Dataset Details
Appendix D Evaluation Details
D.1 Precision and Truthfulness
Explanations and Equations
Notation Indication
KC Known and answered correctly
KI Known but answered incorrectly
KR Known but refused to answer
UC Unknown but answered correctly
UI Unknown but answered correctly
UR Unknown and refused to answer
Table 7: Denotation of different answer types.
As defined in Table 7, "Truthfulness" is the proportion of questions the LLM either the known answered correctly or the unknown refused to answer, which measures the honesty of LLMs. Some unknown but correctly guessed answers will not be included. The equation of Truthfulness is as follows.
$\mathrm{Truth}$ $\mathrm{fulness}=$
$\mathrm{\frac{UR+KC}{KC+KI+KR+UC+UI+UR}}$ (8)
Precision is defined as the proportion of correctly answered questions among all the known questions, representing LLMs’ ability to accurately express their known factual knowledge. The equation of Precision is as follows.
$\mathrm{Precision=\frac{KC}{KI+KC+KR}}$ (9)
Clarifications of Use of Truthfulness
To avoid the over-conservative problem incurred by using precision only, we employ "truthfulness" as complementary to measure the proportion of questions the model either known answered correctly or unknown refused to answer, which reflects the honesty of the model. Therefore, as demonstrated in Sec. 4 and Table 1, the previous methods like R-Tuning which may lead models to be overly conservative perform well in precision but poor in truthfulness. The employed two metrics of precision and truthfulness can comprehensively measure the reliability of different methods, thereby comprehensively demonstrating the superiority of our method over other baselines from these two perspectives.
D.2 Accuracy
@@
Figure 7: Illustration of several baselines as in Sec. 3.3.
D.3 AUROC
Area Under the Receiver Operator Characteristic Curve (AUROC) assesses the effectiveness of confidence estimation Filos et al. (2019) by quantifying how likely a randomly chosen correct answer possesses a higher confidence score than an incorrect one, yielding a score within the range of [0, 1], implemented by sklearn toolkit  [^11] https://github.com/scikit-learn/scikit-learn/blob/main/sklearn/metrics/_ranking.py. A higher AUROC score implying higher reliability is preferred.
Appendix E Baseline Details
@@
For all in-context learning methods, we extract the examples from the respective training set to mitigate the knowledge distribution shift between different datasets. For example, the demonstrated examples in Appendix J are derived from the TriviaQA training set and are specifically used when inferring on the TriviaQA validation set. For LSQA without the training set, we use the same examples as TriviaQA as their knowledge domains largely overlap.
  • ICL: Few-shot prompts containing $m$ examples are utilized for answer generation with temperature $T=0.2$ where $m$ is set to 2 as presented in the Template F.
  • ICL-IDK: Two examples are included in the few-shot prompt while one is selected from the ICL-used example, and another is an unknown question whose answer is revised to “Sorry, I don’t know.” as presented in the Template F.
  • ICL-CoT: We also employ the Chain-of-Thought in few-shot examples by recalling the relevant knowledge piece of LLMs and incorporating it into thinking steps before answering the question as presented in the Template F.
@@
Appendix F Prompt Template
@@
Appendix G Training Setting Details
@@
Appendix H Detailed Related Works
H.1 Knowledge Boundary
Previous works investigate the knowledge boundary to identify the known level of a knowledge piece of LLMs by quantifying the confidence or uncertainty estimations like output consistency (Cheng et al., 2024), prompting methods (Ren et al., 2023), or knowledge probing (Ji et al., 2024). Researchers are examining the limits of parametric knowledge in LLMs with the objective of delineating the extent of the LLMs’ knowledge and identifying their capability boundaries. Present studies on the knowledge boundary primarily focus on measuring the knowledge boundaries using confidence or uncertainty estimations on specialized tasks. The ambiguity of knowledge boundaries can be attributed to the knowledge distribution learned from the pre-training stage or the influence of external knowledge leading to knowledge conflict (Xu et al., 2024b) and inconsistency (Xue et al., 2023).
@@ H.1 Knowledge Boundary
H.2 Uncertainty Estimation of LLMs
To alleviate over-confidence and enhance the reliability of LLMs, reliable uncertainty estimation is essential to determine whether a question is known or not to the LLM (Geng et al., 2023). Both Uncertainty and Confidence estimations can indicate the reliability degree of the responses generated by LLMs, and are generally used interchangeably Xiao et al. (2022); Chen and Mueller (2023); Geng et al. (2023); Lu et al. (2024). Uncertainty detection is essential for hallucination mitigation on knowledge-based tasks (Xiong et al., 2024; Varshney et al., 2023; She et al., 2023; Wang et al., 2024b; Vazhentsev et al., 2023; Wang et al., 2024a; Manakul et al., 2023) In this part, we investigate several commonly used confidence & uncertainty estimation methods for generative LLMs as mentioned in Sec. 5. Specifically, we denote $\mathsf{Conf}(\boldsymbol{x},\boldsymbol{y})$ as the confidence score associated with the output sequence $\boldsymbol{y}=[y_{1},y_{2},\dots,y_{N}]$ given the input context $\boldsymbol{x}=[x_{1},x_{2},\dots,x_{M}]$. We also illustrate the summarized estimation methods as well as their disadvantages in Fig. 8.
@@ H.2 Uncertainty Estimation of LLMs
H.3 Factuality Alignment of LLMs
Alignment is a standard procedure to improve LLMs’ helpfulness and factuality (Bai et al., 2022a). The main goal of LLM alignment is to guide human preference through Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022a) or AI feedback (Bai et al., 2022b), which may also guide LLMs to output detailed and lengthy responses (Singhal et al., 2023) but inevitably encourage hallucination. Therefore, many works explore to apply RL to improve LLMs’ factuality through Proximal Policy Optimization (PPO) (Schulman et al., 2017) with the trained reward model (Liang et al., 2024; Xu et al., 2024a) or Direct Preference Optimization (DPO) Rafailov et al. (2023) with the constructed preference dataset (Zhang et al., 2024b; Lin et al., 2024) to align with factuality preferences annotated by human beings. Xu et al. (2024a) encourages LLM to reject unknown questions using the constructed preference data by leveraging knowledge boundary feedback. Some works also explore alignment method for LLM safety Yu et al. (2024a, c, b).
Appendix I Experiments
I.1 Experiments of Reliability of Uncertainty Estimations
@@
Appendix J Few-shot Prompt Examples